# The Tarantula Massive Binary Monitoring

### v1

@@
Massive O-type stars () with an X-ray quiet black hole (BH) companion represent a crucial stage in massive binary evolution leading to binary BH mergers. The population of such binaries remains elusive, with candidate or confirmed systems. The Tarantula nebula harbors thousands of massive stars, 2–3% of which are expected to have BH companions. It is therefore an ideal place to hunt for such systems. Here we analyse 30 epochs of VLT/FLAMES IFU high-resolution observations of the H region, as well as archival FLAMES spectroscopy, of VFTS 812, a 17-day single-lined spectroscopic binary with an O4 V primary and a minimum secondary mass of . Following careful removal of the nebular contamination, spectral disentangling on the new data did not reveal any signature of the hidden companion. We derive kK, and km s-1 for the O4 V component, yielding a (single star) evolutionary mass of M⊙ and an age in the range of 0-1.6 Myr. Using injection tests of various luminous artificial companions in our data, we exhaustively rule out the presence of any luminous signature from a main sequence (MS) star more massive than . We discuss the possible nature of the companion, suggesting that the rejuvenated O star + BH companion is the most suitable scenario to consistently explain the location, (rejuvenated) young age, eccentricity and lack of companion signature. While this establishes VFTS 812 as a strong candidate O+BH system, follow-up observations are deemed necessary for robust confirmation and to search for accretion signatures on the O4 V star.
@@
Massive binary evolution is a key formation channel for mergers of neutron stars and black holes (BH) that are detected via gravitational waves. The intermediate OB-type star + BH binary stage holds key evolutionary constraints in this channel. Especially when the OB+BH system has not started to interact, its orbital and dynamical properties retain pristine information in the immediate aftermath of core-collapse of the primary (Vigna-Gómez et al., 2024; Willcox et al., 2025). Around 2–3% of OB stars are predicted to host compact companions (Langer et al., 2020). However, known X-ray quiet OB+BH binaries are scarce, with only a handful recently identified (e.g. Mahy et al., 2022; Shenar et al., 2022b). While the Gaia mission is expected to uncover a significant population of OB+BH binaries (Janssens et al., 2022; Nagarajan et al., 2025), spectroscopy will remain the main BH hunting tool for systems with short periods or at large distances such as in the Magellanic Clouds.
@@
VFTS 243 is the only confirmed X-ray quiet OB+BH binary in the Large Magellanic Cloud or LMC (Shenar et al., 2022b), discovered as part of a much broader effort to characterize a sample of 51 O-type single-lined spectroscopic binaries (SB1) in the Tarantula nebula region (Shenar et al., 2022a, referred to as TS22 here onward). A significant fraction of that sample was re-classified as double-lined spectroscopic binaries (SB2) based on spectral disentangling, a technique for the separation of component spectra in spectroscopic binaries (e.g. Hadrava, 1995). Others retained their SB1 status or were flagged as uncertain due to nebular contamination impeding the disentangling process. Subsequently, an improved observational approach was deemed necessary to better identify OB+BH systems in the sample.
@@
VFTS 812, reported as a 17-day eccentric () binary with an O4 V primary weighing about and a minimum secondary mass of , was one such system classified as uncertain (TS22). A BH companion to the O4 V primary would likely descend from a similarly or more massive progenitor, and could potentially be massive itself. Given the minimum secondary mass is in the BH mass regime, we revisit VFTS 812 with new observations aiming to perform better disentangling and establish the nature of its companion. In this letter, we present the observations, data reduction, disentangling analysis and possible nature of the unseen companion.
2 Observations, data reduction and sky subtraction
@@
Prior data for VFTS 812 were obtained with the GIRAFFE spectrograph fed with single FLAMES-MEDUSA fibers on the Very Large Telescope (VLT) in the low resolution modes LR02 and LR03, along with some additional high resolution data around the Hα region (Evans et al., 2011; Almeida et al., 2017). In this work, VFTS 812 was observed with the Integral Field Unit (IFU) fiber mode of the FLAMES GIRAFFE spectrograph with the HR03 grating (4032 – 4203 , ). This setting was motivated by two expected improvements: (i) a better nebular subtraction compared to single-fiber observations; (ii) higher resolution enabling more robust spectral disentangling.
@@
Between October 2023 and September 2025, a total of 30 epochs were obtained under ESO programs 112.260M and 114.275G (see Appendix A). Raw data were first corrected for cosmic rays using Astro-SCRAPPY111github.com/astropy/astroscrappy (van Dokkum, 2001). We then used the FLAMES-IFU data reduction recipes within Esorex version 3.13.7 to reduce all our data. Each IFU comprises 20 individual fibers bundled in a configuration barring the corners, where each fiber (referred to as spaxel here onward) records its own spectrum. Under ideal conditions, four spaxels at the center of the IFU should cover the star, whereas remaining spaxels can be used to obtain the sky background in its immediate vicinity. Poor seeing or IFU misalignment are two main factors that can adversely affect the data quality.
@@
Despite measures taken to improve sky subtraction (see Appendix B), it was non-trivial to choose appropriate star and sky spaxels for perfect subtraction. While there are clear improvements compared to the low resolution single-fiber data, the spatial as well as temporal variability of the nebular line made it challenging to have no residuals after subtraction, and subsequently to obtain spectra appropriate for spectral disentangling. We circumvent this by implementing error spectra in our disentangling approach which can account for imperfect subtraction (see Section 4). A detailed description of our optimal sky subtraction strategy is discussed in Appendix B. In total, spectra for 26 out of 30 epochs were successfully extracted and used for analysis in this work (Table 1).
@@ 3 Constraining component masses
For an SB1 system with a known mass function, an accurate estimate of the primary mass can place strong constraints on the minimum mass of the secondary. VFTS 812 has a mass function and a primary spectral type of O4 V (TS22). The system does not show eclipses in its light curve, making dynamical masses inaccessible without an astrometric orbit. Previous estimates for VFTS 812 heavily relied on approximate evolutionary masses, with TS22 reporting derived by taking the mean evolutionary mass of all single stars of similar spectral type from the 30 Doradus sample of Schneider et al. (2018). Here, we revisit the primary mass estimates using detailed atmospheric analysis and subsequent constraints on the secondary mass.
@@ 3.1 Spectroscopic mass of the primary
We compiled all archival and new FLAMES data for VFTS 812, obtained with the gratings LR02 (397–457 nm), LR03 (450–508 nm), HR03 (403–420 nm) and HR15N (647–679 nm). The spectra from each of the gratings were normalized and coadded in the reference frame of the primary star, clipping out the nebular emission lines before coaddition. We fit fastwind atmospheric models (Santolaya-Rey et al., 1997; Sundqvist and Puls, 2018) on the combined spectrum to obtain stellar parameters such as effective temperature (), surface gravity (log ) and helium abundance (), among others (see full list in Table 3). The fitting was done with Kiwi-GA, which explores the stellar parameter space allowing for the calculation of statistically derived error margins (Brands et al., 2022, 2025; Verhamme et al., 2024). More details on spectral fitting are discussed in Appendix C.
Because of a lack of He i and nitrogen lines, the derived temperature has large error-margins. The absence of both N v and N iv lines results in only an upper limit on the nitrogen abundance. The derived spectroscopic mass of the star is , slightly lower than the evolutionary mass reported in TS22. This is not fully unexpected as spectroscopic masses are often lower than the derived evolutionary masses (Herrero et al., 1992). This is potentially due to the missing turbulent pressure in spectral codes such as fastwind, which can be significant (Moens et al., 2025). Accounting for this turbulence results in an increase in log , which can increase the mass significantly (González-Torà et al., 2025, Verhamme et al. in prep.). Moens et al. (2025) found a typical velocity corresponding to the turbulent pressure for an O4 V star to be around 50 . This velocity does not affect the equivalent width of the lines like the micro-turbulent velocity would and is physically more akin to the macro-turbulent velocity. However, this velocity is not used in the broadening of the line profiles in the formal solution. If we include this turbulent pressure in fastwind, we find an increased spectroscopic mass of , stemming primarily from the increased log value while other parameters remain consistent (Table 3). Since choosing different values will result in significantly different masses, we adopt the parameters derived without including turbulent pressure and turn to evolutionary masses as the more suitable estimate for VFTS 812.
@@ 3.2 Evolutionary mass of the primary
With newly performed atmospheric analysis, we obtained an effective temperature and bolometric luminosity log . The primary star was not found to have any significant helium or nitrogen enrichment that could indicate past interactions with the companion. We used the BONNSAI Bayesian tool (Schneider et al., 2014) to fit single star stellar tracks from Brott et al. (2011) to the primary using and log as inputs. Surface gravity was not included given its strong dependence on turbulent pressure which is unconstrained.
BONNSAI successfully replicated all observables consistently within 1-, although with a low mass tail of high projected rotational velocity solutions. Given the modest of km s-1 measured from atmospheric analysis, we excluded all BONNSAI solutions beyond a rotational velocity of 300 km s-1. Subsequently, we constrained the evolutionary mass of the O4 V star to , the radius , and an estimated age of Myr (all values are modes). This mass is significantly higher than the corresponding spectroscopic mass (). Additionally, the evolutionary mass derived using the values with turbulent pressure included is essentially identical ().
@@ 3.3 Secondary mass and photometric constraints
We combine the orbital parameters of VFTS 812 from TS22 with the mass and radius posterior distributions obtained from BONNSAI to obtain a distribution for the minimum secondary mass which gives at 5% significance. Assuming an isotropic orientation of the orbit, we obtain a secondary mass distribution with at 5% significance.
VFTS 812 shows no significant photometric variability or eclipses in OGLE observations (TS22). The absence of eclipses can put an upper limit on the inclination of the binary, assuming that the faint secondary is large enough to eclipse the bright primary. Assuming an MS companion and a mass-radius relation obtained from zero-age main sequence stars in the BONNSAI catalog, we use the absence of a primary eclipse to further constrain the value of to exceed at 5% significance. More on photometric constraints is discussed in Appendix E.
@@ 4 Spectral disentangling
As discussed in Section 2, the individual epoch spectra of VFTS 812 still have residual features from nebular line subtraction. To minimize the effects of these residuals on the disentangling, we implemented wavelength-dependent error propagation in the shift-and-add technique of TS22222github.com/TomerShenar/Disentangling_Shift_And_Add (see their Section 3.1). We used an error spectrum for each input science spectrum, making it possible to assign negligible weight to the region contaminated by the nebular line. For the HR03 setup, the wavelength region from 4103.9 – 4106.7 was most affected by the nebular line, and thus essentially excluded from the disentangling process. A mathematical description of our modification to the shift-and-technique of TS22 is described in Appendix D.
@@ 4.1 Searching for a companion
We adopted the orbital solution for VFTS 812 reported by Almeida et al. (2017), with an updated time of periastron passage derived from new observations (Table 4). Taking an agnostic approach, we performed two-component disentangling over a grid of component radial velocity semi-amplitudes ranging from 30 – 52 km s-1 in steps of 1 km s-1 and 10 – 500 km s-1 in steps of 2.5 km s-1, respectively. This ensured a thorough exploration of different mass ratios. Figure 1 shows a 2-D chi-squared map of our disentangling analysis. The primary semi-amplitude was recovered at km s-1, consistent with km s-1 reported in TS22. , however, was unconstrained as evident from the chi-squared map. The disentangled spectrum of the primary is consistent with the O4 V primary, while that for the secondary is essentially flat and featureless (Fig. 2). The latter translates to an absence of a detectable companion, either because it is too faint or dark. To further investigate which companion types can avoid detection in our spectra, we carried out an injection study, as described in the following subsection.
@@ 4.2 Limits on the flux and mass of the secondary
To test our ability of detecting companions of different masses, we performed an injection study wherein we injected companion model spectra into the data and attempted to retrieve them with spectral disentangling. Given the O4 V primary and assuming the secondary is an MS star, we used BONNSAI to determine the radius, temperature and luminosity of the injected companion by fixing its age to that of the primary. We computed fastwind model spectra based on these inputs, and convolved them with the Bessel B filter to estimate their flux contribution (Table 5). Subsequently, we synthesized mock spectra by injecting an MS secondary in the data using the computed flux contribution and with an RV semi-amplitude corresponding to its mass. Mock datasets were made for MS companions of 5, 5.5, 6, 6.5, 7, 8 and 10 . For every dataset, we ran a companion search using disentangling. The disentangled secondary spectrum was retrieved at the same value that it was injected at, while keeping the grid same as Section 4.1. Since the real is not known, we retrieved the secondary spectrum at different values for the no-injection case to search for spectral features.
The results for select cases are summarized in Figure 2, along with the companion search run from Section 4.1 with no injected secondary. Each panel shows the disentangled secondary compared with the injected model (flat for the no-injection case). We performed an -test in the 4072–4130 range to determine whether the model is a better fit to the retrieved spectrum than a featureless flat line, for which we used a 5% significance threshold. The continuum region was determined outside the 4090–4112 region and fixed for both the flat line and model fits. Subsequently, chi-squared values were evaluated in the full 4072–4130 range to compare the fits.
Assuming rotational velocities of 0, 250 and 500 km s-1 for the injected models, we are able to successfully retrieve companions with masses 6 and above. For the no-injection case, the flat line vs model fits are statistically indistinguishable when tested with 5, 5.5 and 6 models. We can therefore confidently rule out a luminous MS companion.

### v2

@@
Massive O-type stars () with an X-ray-quiet black hole (BH) companion represent a crucial stage in the massive binary evolution leading to binary BH mergers. The population of such binaries remains elusive, with candidate or confirmed systems. The Tarantula nebula harbors thousands of massive stars, 2–3% of which are expected to have BH companions. It is therefore an ideal place to hunt for such systems. We analyzed 30 epochs of VLT/FLAMES IFU high-resolution observations of the H region and archival FLAMES spectroscopy of VFTS 812, a 17-day single-lined spectroscopic binary with an O4 V primary and a minimum secondary mass of . Following careful removal of the nebular contamination, spectral disentangling on the new data did not reveal any signature of the hidden companion. We derive kK, and km s-1 for the O4 V component, yielding a (single-star) evolutionary mass of M⊙ and an age in the range 0-1.6 Myr. Using injection tests for various luminous artificial companions in our data, we exhaustively ruled out the presence of any luminous signature from a main sequence star more massive than . We discuss the possible nature of the companion, suggesting that a rejuvenated O star + BH companion is the most suitable scenario to consistently explain the location, (rejuvenated) young age, eccentricity, and lack of companion signature. While this establishes VFTS 812 as a strong candidate O+BH system, follow-up observations are deemed necessary for a robust confirmation and to search for accretion signatures on the O4 V star.
@@
Massive binary evolution is a key formation channel for the mergers of neutron stars and black holes (BHs) that are detected via gravitational waves. Knowledge of the intermediate OB-type star + BH binary stage would allow us to place key constraints on this evolutionary channel. In the immediate aftermath of the core-collapse of the primary, especially when the OB+BH system has not started to interact, its orbital and dynamical properties retain pristine information (Vigna-Gómez et al., 2024; Willcox et al., 2025). Around 2–3% of OB stars are predicted to host compact companions (Langer et al., 2020). However, known X-ray-quiet OB+BH binaries are scarce, with only a handful recently identified (e.g., Mahy et al., 2022; Shenar et al., 2022b). While the Gaia mission is expected to uncover a significant population of OB+BH binaries (Janssens et al., 2022; Nagarajan et al., 2025), spectroscopy will remain the main BH hunting tool for systems with short periods or at large distances, such as in the Magellanic Clouds.
@@
VFTS 243 is the only confirmed X-ray-quiet OB+BH binary in the Large Magellanic Cloud (Shenar et al. 2022b). It was discovered as part of a much broader effort to characterize a sample of 51 O-type single-lined spectroscopic binaries (SB1) in the Tarantula nebula region (Shenar et al., 2022a, TS22 hereafter). A significant fraction of that sample was reclassified as double-lined spectroscopic binaries (SB2) based on spectral disentangling, a technique for the separation of component spectra in spectroscopic binaries (e.g., Hadrava, 1995). Others retained their SB1 status or were flagged as uncertain due to nebular contamination impeding the disentangling process. Subsequently, an improved observational approach was deemed necessary to better identify OB+BH systems in the sample.
@@
VFTS 812, reported as a 17-day eccentric () binary with an O4 V primary weighing about and a minimum secondary mass of , was one such system classified as uncertain (TS22). A BH companion to the O4 V primary would likely descend from a similarly or more massive progenitor, and could potentially be massive itself. Given that the minimum secondary mass is in the BH mass regime, we revisited VFTS 812 with new observations aiming to perform a better disentangling analysis and establish the nature of its companion. In this letter we present the observations, data reduction, disentangling analysis, and possible nature of the unseen companion.
2 Observations and data reduction
@@
Prior data for VFTS 812 were obtained with the GIRAFFE spectrograph fed with single FLAMES-MEDUSA fibers on the Very Large Telescope (VLT) in the low-resolution modes LR02 and LR03, along with some additional high-resolution data around the Hα region (Evans et al., 2011; Almeida et al., 2017). In this work, VFTS 812 was observed with the integral field unit (IFU) fiber mode of the GIRAFFE spectrograph with the HR03 grating (4032 – 4203 , ). The choice of this setting was motivated by two expected improvements: (i) a better nebular subtraction compared to single-fiber observations and (ii) a higher resolution enabling more robust spectral disentangling.
@@
Between October 2023 and September 2025, a total of 30 epochs were obtained under ESO programs 112.260M and 114.275G (see Appendix B). Raw data were first corrected for cosmic rays using Astro-SCRAPPY111github.com/astropy/astroscrappy (van Dokkum, 2001). We then used the FLAMES-IFU data reduction recipes within Esorex version 3.13.7 to reduce all our data. Each IFU comprises 20 individual fibers bundled in a configuration barring the corners, where each fiber (referred to as a ”spaxel” hereafter) records its own spectrum. Under ideal conditions, four spaxels at the center of the IFU should cover the star, and the remaining spaxels can be used to obtain the sky background in its immediate vicinity. Poor seeing or IFU misalignment are two main factors that can adversely affect the data quality.
@@
Despite measures taken to improve sky subtraction (see Appendix C), it was nontrivial to choose appropriate star and sky spaxels for perfect subtraction. While there are clear improvements compared to the low-resolution single-fiber data, the spatial as well as temporal variability of the nebular line made it challenging to have no residuals after subtraction, and subsequently to obtain spectra appropriate for spectral disentangling. We circumvented this by implementing error spectra in our disentangling approach, which can compensate for imperfect subtraction (see Sect. 4). A detailed description of our optimal sky subtraction strategy is discussed in Appendix C. In total, spectra for 26 out of 30 epochs were successfully extracted and used for analysis in this work (Table 1.1).
@@ 3 Constraining component masses
For an SB1 system with a known mass function, an accurate estimate of the primary mass can place strong constraints on the minimum mass of the secondary. VFTS 812 has a mass function and a primary spectral type of O4 V (TS22). The system does not show eclipses in its light curve, making it impossible to determine dynamical masses without an astrometric orbit. Previous estimates for VFTS 812 heavily relied on approximate evolutionary masses; TS22 report , which they derived by taking the mean evolutionary mass of all single stars of a similar spectral type from the 30 Doradus sample of Schneider et al. (2018). We revisited the primary mass estimates using detailed atmospheric analysis and subsequent constraints on the secondary mass.
@@ 3.1 Spectroscopic mass of the primary
We compiled all archival and new FLAMES data for VFTS 812, obtained with the gratings LR02 (397–457 nm), LR03 (450–508 nm), HR03 (403–420 nm), and HR15N (647–679 nm). The spectra from each of the gratings were normalized and co-added in the reference frame of the primary star; we clipped out the nebular emission lines before the co-addition. We fit fastwind atmospheric models (Santolaya-Rey et al., 1997; Sundqvist and Puls, 2018) to the combined spectrum to obtain stellar parameters such as the effective temperature (), surface gravity (log ), and helium abundance (), among others (see the full list in Table 3). The fitting was done with Kiwi-GA, which explores the stellar parameter space, thereby enabling the calculation of statistically derived error margins (Brands et al., 2022, 2025; Verhamme et al., 2024). More details on spectral fitting are discussed in Appendix D.
Because of a lack of He i and nitrogen lines, the derived temperature has large error margins. The absence of both N v and N iv lines means we can only place an upper limit on the nitrogen abundance. The derived spectroscopic mass of the star is , slightly lower than the evolutionary mass reported in TS22. This is not fully unexpected as spectroscopic masses are often lower than the derived evolutionary masses (Herrero et al., 1992). This is potentially due to the missing turbulent pressure in spectral codes such as fastwind, which can be significant (Moens et al., 2025). Accounting for this turbulence results in an increase in log , which can increase the mass significantly (González-Torà et al., 2025, Verhamme et al. in prep.). Moens et al. (2025) found a typical velocity corresponding to the turbulent pressure for an O4 V star to be around 50 . This velocity does not affect the equivalent width of the lines like the micro-turbulent velocity would and is physically more akin to the macro-turbulent velocity. However, this velocity is not used in the broadening of the line profiles in the formal solution. If we include this turbulent pressure in fastwind, we find an increased spectroscopic mass of ; this stems primarily from the increased log value (the other parameters remain consistent; Table 3). Since choosing different values will result in significantly different masses, we adopted the parameters derived without including turbulent pressure and considered the evolutionary masses to be the more suitable estimate for VFTS 812.
@@ 3.2 Evolutionary mass of the primary
With this new atmospheric analysis, we obtained an effective temperature and bolometric luminosity log . The primary star was not found to have any significant helium or nitrogen enrichment that could indicate past interactions with the companion. We used the BONNSAI Bayesian tool (Schneider et al., 2014) to fit single-star stellar tracks from Brott et al. (2011) to the primary using and log as inputs. Surface gravity was not included given its strong dependence on turbulent pressure, which is unconstrained.
BONNSAI successfully replicated all observables consistently within 1, although with a low mass tail of high projected rotational velocity solutions. Given the modest of km s-1 measured from the atmospheric analysis, we excluded all BONNSAI solutions beyond a rotational velocity of 300 km s-1. Subsequently, we constrained the evolutionary mass of the O4 V star to , the radius , and estimated an age of Myr (all values are modes). This mass is significantly higher than the corresponding spectroscopic mass (). Additionally, the evolutionary mass derived using the values with turbulent pressure included is essentially identical ().
@@ 3.3 Secondary mass and photometric constraints
We combined the orbital parameters of VFTS 812 from TS22 with the mass and radius posterior distributions obtained from BONNSAI to obtain a distribution for the minimum secondary mass, obtaining at 5% significance. Assuming an isotropic orientation of the orbit, we obtained a secondary mass distribution with at 5% significance.
VFTS 812 shows no significant photometric variability or eclipses in OGLE observations (TS22). The absence of eclipses means we can put an upper limit on the inclination of the binary, assuming that the faint secondary is large enough to eclipse the bright primary. Assuming a main sequence (MS) companion and a mass–radius relation obtained from zero-age MS stars in the BONNSAI catalog, we used the absence of a primary eclipse to further constrain the value of to exceed at 5% significance. Photometric constraints are discussed further in Appendix E.
@@ 4 Spectral disentangling
As discussed in Sect. 2, the individual epoch spectra of VFTS 812 still have residual features from nebular line subtraction. To minimize the effects of these residuals on the disentangling, we implemented wavelength-dependent error propagation in the shift-and-add technique of TS22222github.com/TomerShenar/Disentangling_Shift_And_Add (see their Sect. 3.1). We used an error spectrum for each input science spectrum, which makes it possible to assign negligible weight to the region contaminated by the nebular line. For the HR03 setup, the wavelength region from 4103.9 – 4106.7 was most affected by the nebular line and was thus essentially excluded from the disentangling process. A mathematical description of our modification to the shift-and-add technique of TS22 is presented in Appendix F.
@@ 4.1 Searching for a companion
We adopted the orbital solution for VFTS 812 reported by Almeida et al. (2017), with an updated time of periastron passage derived from new observations (Table 4). Taking an agnostic approach, we performed two-component disentangling over a grid of component radial velocity (RV) semi-amplitudes, and , which range from 30 – 52 km s-1 in steps of 1 km s-1 and 10 – 500 km s-1 in steps of 2.5 km s-1, respectively. This ensured a thorough exploration of different mass ratios. Figure 1 shows a 2D chi-squared map of our disentangling analysis. The primary semi-amplitude, , was recovered at km s-1, consistent with the km s-1 reported in TS22. , however, was unconstrained, as is evident from the chi-squared map. The disentangled spectrum of the primary is consistent with the O4 V primary, while that for the secondary is essentially flat and featureless (Fig. 2). The latter translates to an absence of a detectable companion, either because it is too faint or dark. To further investigate which companion types can avoid detection in our spectra, we carried out an injection study, as described in the following subsection.
@@ 4.2 Limits on the flux and mass of the secondary
To test our ability to detect companions of different masses, we performed an injection study wherein we injected companion model spectra into the data and attempted to retrieve them via spectral disentangling. Given the O4 V primary and assuming the secondary is an MS star, we used BONNSAI to determine the radius, temperature, and luminosity of the injected companion by fixing its age to that of the primary. We computed fastwind model spectra based on these inputs, and convolved them with the Bessel B filter to estimate their flux contribution (Table 5). Subsequently, we synthesized mock spectra by injecting an MS secondary into the data using the computed flux contribution and with an RV semi-amplitude corresponding to its mass. Mock datasets were made for MS companions of 5, 5.5, 6, 6.5, 7, 8, and 10 . For every dataset, we ran a companion search using disentangling. The disentangled secondary spectrum was retrieved at the same value that it was injected at while keeping the grid the same as presented in Sect. 4.1. Since the real is not known, we retrieved the secondary spectrum at different values for the no-injection case to search for spectral features.
The results for select cases are summarized in Fig. 2, along with the companion search run from Sect. 4.1 with no injected secondary. Each panel shows the disentangled secondary compared with the injected model (flat for the no-injection case). We performed an -test in the 4072–4130 range to determine whether the model is a better fit to the retrieved spectrum than a featureless flat line, for which we used a 5% significance threshold. The continuum region was determined outside the 4090–4112 region and fixed for both the flat line and model fits. Subsequently, chi-squared values were evaluated in the full 4072–4130 range to compare the fits.
Assuming rotational velocities of 0, 250, and 500 km s-1 for the injected models, we were able to successfully retrieve companions with masses of 6 and above. For the no-injection case, the flat line versus model fits are statistically indistinguishable when tested with 5, 5.5, and 6 models. We can therefore confidently rule out a luminous MS companion.